4. Suppose that Dr. Heart has his doctors office in the same building as Mr. Grahams bakery. Unfortunately, for Dr. Heart,
Answer #1

(A) For finding out the socially efficient option, we can compare the total profits for both individuals in each case:
1. Noisy Machinery: Total Profits= 175000 + 100000 = 275,000
2. Quiet Machinery: Total Profits= 200000 + 80000 = 280,000
Hence, clearly, quiet machinery is socially efficient.

(B) If the two cannot negotiate, Mr Graham will use the noisy machinery as it yields higher profits for him, he will only be concerned with his own profits

(C) If the two negotiate, Mr Graham could be made to use quieter machinery by compensating him for the profit that he loses by using the quieter machinery.
As he loses 20,000 by using quieter machinery, Dr Heart could offer him 20,000 as compensation, after which he will be indifferent to the machinery being used. At the same time, Dr Heart''s profits would increase by 25,000, and after compensation, he still ends up with additional profits of 5,000.
So, this situation would be beneficial for both!

Hope it helps